OpenStudy (anonymous):

how do we do addition method with 3 equations? 4x+8y-4z=4 3x+8y+5z=-11 -2x+y+12z=-17

OpenStudy (anonymous):

Which three equations? Write them down.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

4x+8y-4z=4 3x+8y+5z=-11 -2x+y+12z=-17

OpenStudy (anonymous):

The addition method is also called the elimination method. Do you know how to do the elimination method ?

OpenStudy (anonymous):

yes, i got up to where i got -x+4z=-15, but i dont know how to do the other steps

OpenStudy (anonymous):

ok...I see you cancelled out the y's....let me show you...hold on...elimination methodisnot hard, just time consuming

OpenStudy (anonymous):

okay :]

OpenStudy (anonymous):

4x + 8y - 4z = 4 --> (-1)4x + 8y - 4z = 4 3x + 8y + 5z = -11 ----------------- -4x - 8y + 4z = -4 3x + 8y + 5z = - 11 ------------------add -x + 9z = - 15 Is this how you started....because I got something different.

OpenStudy (anonymous):

oh yeah thats what i got , i just typed it wrong

OpenStudy (anonymous):

now you have to take the last to equations and do the same thing....just remember, what you cancelled out in the first two equations, you have to cancel out in the last two equations. So, do the last two equations and cancel out the y's as well. Try that and tell me what you get

OpenStudy (anonymous):

oops....the last two equations....

OpenStudy (anonymous):

wait, i do 3x+8y+5z=-11 -2x+y+12z=-17 ?

OpenStudy (anonymous):

okay...you have to find a way to cancel out the y's. Try multiplying the last equation by a -8.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay i got 19x-91z=-125

OpenStudy (anonymous):

3x + 8y + 5z = -11 -2x + y + 12z = - 17 -->(-8)-2x + y + 12z = - 17 ----------------- 3x + 8y + 5z = - 11 16x - 8y - 96z = 136 (result of multiplying by -8) ------------------add 19x - 91z = 125 almost.... now you take the answers you got...the two that you eliminated the y's from, add them and eliminate any one of those variables to find the other variable. Do you understand ?

OpenStudy (anonymous):

oh okay, so we should try to cancel out the same variables to get the two equations?

OpenStudy (anonymous):

yes.....and now that you have the two equations, the ones where you cancelled out the y's....do the same thing with them and cancel out a variable. Do you understand or do you want me to show you ?

OpenStudy (anonymous):

i understand , but i want to make sure i did it right

OpenStudy (anonymous):

I will show you..... these are the two equations you have where the y's have been cancelled out : -x +9z = - 15 -->(19)-x + 9z = - 15 19x - 91z = 125 --------------- -19x + 171z = - 285 (result of multiplying by 19) 19x - 91z = 125 ---------------add 0 + 80z = - 160 x = - 160/80 x = - 2 do you understand what I just did ?

OpenStudy (anonymous):

isnt it 60? and z?

OpenStudy (anonymous):

my bad....it is z....but it is not 60. How did you get 60 ?

OpenStudy (anonymous):

wait, i messed up , i got 80 too

OpenStudy (anonymous):

good. Now you can substitute the -2 in for z in either of the equations that the y was cancelled out of . Do you understand ?

OpenStudy (anonymous):

yeah, so do i get x=3

OpenStudy (anonymous):

how would i get y tho?

OpenStudy (anonymous):

let me check... -x +9z = -15 -x + 9(-2) = - 15 -x - 18 = - 15 -18 + 15 = x -3 = x or you could have done this one... 19x - 91z = 125 19x - 91(-2) = 125 19x + 182 = 125 19x = 125 - 182 19x = - 57 x = - 3 either way you will get x = - 3 do you understand this ?

OpenStudy (anonymous):

yeah, i understand that part but what about y?

OpenStudy (anonymous):

To get y, you just plug in your known variables into any of the three original equations and you should be able to find y.Once you know all 3 variables, you can check by subbing in all known variables into any of the original equations and if they come out equal, your answers are correct.

OpenStudy (anonymous):

oh okay, i get it :] thank you very much!

OpenStudy (anonymous):

you are very welcome. Be careful with elimination method because it is easy to make mistakes. And if you make a mistake at the beginning, it will travel through the rest of the problem and you won't get the right answer .

OpenStudy (anonymous):

oh okay , yeah, i had messed up and i noticed it did mess up the problem . :]

OpenStudy (anonymous):

so did I....I told you x when it should have been z...lol

OpenStudy (anonymous):

lol its okay, you helped me out a lot! thank you very much once again, i appreciate .

OpenStudy (anonymous):

it.

OpenStudy (anonymous):

anytime :)